NPP Projection Admin 1/30
To do this evening/ was completed: Still waiting on Monday's prints but they should arrive this evening.
2/6 playlists in progress. Forecasts ingested. Old playlists removed from servers.
Shrinking Man tech'd, volume logged.
Blazing Saddles blu-ray delivered, ripped, DCP going in T1.
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content):
Outgoing/Shipping (content):
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: I replaced the bulb in Theater 1 due to the flicker. I decided to keep the one that was in the projector for now, it hadn't reached warranty hours yet. Might use it as a backup if it works, we'll see. -AL
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content):
Outgoing/Shipping (content):
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: I replaced the bulb in Theater 1 due to the flicker. I decided to keep the one that was in the projector for now, it hadn't reached warranty hours yet. Might use it as a backup if it works, we'll see. -AL
Theater 2 projector channels had reverted to the wrong measured color files, which is why it had a magenta tint. I changed them back and it looks much better now. -AL
The 2 open electrical boxes in the Theater 2 booth are now covered. -AL
Cue sheets updated/checked (every shift):
Marquee updated?:
Posters updated?: Took down Housemaid and replaced it with Arco -AL
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Cue sheets updated/checked (every shift):
Marquee updated?:
Posters updated?: Took down Housemaid and replaced it with Arco -AL
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment